OpenStudy (anonymous):

2(x + 6) − 10 = 12 + 4(x − 1 (please explain ) i hope you see this mathstudent55

OpenStudy (dls):

@mathstudent55

OpenStudy (anonymous):

i didnt know you can tag somebody i like how that person explains it though

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

2(x + 6) − 10 = 12 + 4(x − 1) You should first distribute 2(x + 6) and 4(x − 1) Do you know how do the distributive property?

OpenStudy (anonymous):

yea so it wwould be 2x+12=12+4x-4 ?

OpenStudy (anonymous):

yeah! ^^ 2(x + 6) − 10 = 12 + 4(x − 1) 2x+12-10=12+4x-4 Now you combine like terms to make the problem easier to solve. Do you know how to combine like terms?

OpenStudy (anonymous):

major issues on that

OpenStudy (anonymous):

Okay, no worries. When i first heard of it i panicked. Combing like terms is just adding,or subtracting numbers that have the same variable or exponent. ie: 3x+5xy-x It helps me when I circle the numbers with the same variable or exponents and the operation in front of it, and do the same with the other numbers except with differ shapes. Because it helps me to distinguish the variables and helps me know which operation to use. 3x+5xy-x combine like terms becomes 2x+5xy do you get how to combine like terms now?

OpenStudy (anonymous):

is that a diffrent example or from the problem show me from the prob >.<

OpenStudy (anonymous):

Differ. :3

OpenStudy (anonymous):

O.O it would help if you showed me >.< step by step!!!! ahhh!!! >.<

OpenStudy (anonymous):

Ah. Sorry. 2(x + 6) − 10 = 12 + 4(x − 1) Distribute 2x+12-10=12+4x-4 Combine like terms 2x+2=8+4x Move all the x's to one side...smaller to bigger 2x+2=8+4x -2x -2x ------------ 2=8+2x solve for x 2=8+2x -8 -8 ----------- -6=2x -- -- 2 2 --------- x= ?

OpenStudy (jhannybean):

Hope you understand my methods :D

OpenStudy (anonymous):

huh? does it equal x+-2 ? im confused what did you get wrong???

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

@KawasumiKimiko does it equal -3

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

thanks

OpenStudy (jhannybean):

And then you plug x=-3 back in to your equation to check to see if it is a correct solution. 2(x+6)−10=12+4(x−1) when x=-3 2(-3+6) -10 = 12 + 4(-3-1) 2(3)-10 = 12 +4(-4) 6 -10 = 12 - 16 -4 = -4 \(\checkmark\)